Input the equation of the given points in standard form not slope intercept form.
(-3,2) (0,3)

Answer by mananth(16946)

x1 y1 x2 y2
-3 2 0 3

slope m = (y2-y1)/(x2-x1)
( 3 - 2 )/( 0 - -3 )
( 1 / 3 )
m= 1/3
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( -3 , 2 ) in
Y = m x + b OK
2.00 = 1/3 * -3 + b
2.00 = -1 + b
b= 2 + 3
b= -1
So the equation will be
Y = 1/3 x + 3
Or x- 3y = -9
m.ananth@hotmail.ca